# Unveiling quantum phase transitions from traps in variational quantum algorithms

### v1

@@ III.3 Extended Su–Schrieffer–Heeger Model
In scenarios where the anisotropy parameter $\delta$ is small, increasing the coupling strength $g$ from 0 to 1 drives a phase transition from a trivial to a topological phase within the system [45]. Initially, the trivial phase displays a dimerized configuration, with spins forming singlet pairs predominantly with their nearest neighbours. As $g$ approaches 1, the system evolves into a topological phase, marked by the emergence of edge states that represent localized excitations at the system’s boundaries. Conversely, when $\delta$ exceeds a critical threshold (approximately $\delta^{\star}=1.6$), the system’s phase diagram becomes more complex, maintaining the trivial and topological phases while also exhibiting a symmetry-broken phase. In this new phase, the increased anisotropy leads to spontaneous symmetry breaking characterized by an antiferromagnetic order with alternating spin alignment.
@@
    $U^{\text{triv}}_{\text{eSSH}}(\bm{\beta},\bm{\gamma})=\prod_{j=1}^{p}$ $\exp\left(-i\frac{\beta_{j}}{2}\mathcal{H}_{xy}\right)\exp\left(-i\frac{\gamma_{j}}{2}\mathcal{H}_{zz}\right)$ (31)
@@
    $U^{\text{topo}}_{\text{eSSH}}(\bm{\beta},\bm{\gamma})=\prod_{j=1}^{p}$ $\exp\left(-i\frac{\beta^{\prime}_{j}}{2}\mathcal{H}^{\prime}_{xy}\right)\exp\left(-i\frac{\gamma^{\prime}_{j}}{2}\mathcal{H}^{\prime}_{zz}\right)$ (33)
@@
The optimization grid for the trivial phase initialization spans from $g_{\text{min}}=0$ to $g_{\text{max}}=0.6$, with a sampling resolution of $\delta g=0.05$. Conversely, the grid for the topological phase initialization ranges from $g_{\text{min}}=0.4$ to $g_{\text{max}}=1$, with the same sampling resolution of $\delta g=0.05$. Given the emphasis on symmetry-oriented phase transitions within this investigation, $\mathcal{P}_{f}$ encompasses all non-identity, reflection symmetric Pauli operators on subsystem $I$, detailed as follows:
@@
The simultaneous acquisition of their expectation values is facilitated through the execution of joint Bell measurements on pairs of qubits within $I$ that are symmetrically positioned, such as $(Q_{n/4},Q_{3n/4-1})$, $(Q_{n/4+1},Q_{3n/4-2})$, and so forth, up to $(Q_{n/2-1},Q_{n/2})$. Fig. 7(a) displays these joint Bell measurements performed on the corresponding qubit pairs in subsystems $I_{1}$ and $I_{2}$, illustrating the method for deriving the feature vector for a 16-qubit quantum chain.
@@
Figure 7: Overview of phase transition detection for the extended Su-Schrieffer-Heeger (eSSH) model. (a) Illustration of the eSSH model’s interaction scheme and the procedure for conducting joint Bell measurements on symmetrically positioned qubit pairs across subsystems $I_{1}$ and $I_{2}$ to derive the feature vector. (b) Fidelity comparison between the optimized quantum states and the true ground states, evaluated across various $g$ and $\delta$ values. (c) Plot of both the original and Gaussian-filter-smoothed loss landscapes for various $\delta$ values, underscoring the identification of critical points. Note that for $\delta=2,3,4$ two minima can be identified, signalling the presence of the two possible phase transitions between the trivial and topological phases, and between the trivial and symmetry-broken ones. (d) Comparison between the predicted phase boundaries, represented by green and blue hollow triangles (indicative of transitions computed from optimized and ground states using $\widetilde{\mathcal{Z}}_{\mathcal{R}}$), and solid black diamonds representing the phase boundaries predicted by the Transformer model. This comparison is set against the phase diagram from Ref. [22], derived using the partial-reflection many-body topological invariant $\widetilde{\mathcal{Z}}_{\mathcal{R}}$. (e) Application of learned order parameters to the modified eSSH model with a transverse field ($\mathcal{H}_{\text{T-eSSH}}$), where $\delta=4$. This panel tracks the phase transitions as the field strength $h$ varies, with phase transitions identified where the Transformer model predicts shifts in phase labels, showcasing the adaptability of the learned parameters across different physical scenarios.
@@
To implement our algorithm, we employ a Set Transformer-based regressor, described in Section II.2. This model is configured with four attention heads and operates at a learning rate of 0.001. Initially, the model applies a self-attention mechanism to the input features to discern dependencies and relationships. Following this, the architecture incorporates a predictive layer that performs a linear transformation, mapping the input features to an intermediate vector of dimension 128. This transformation is augmented by the ReLU activation function, which introduces non-linearity, and is followed by another linear transformation that reduces the feature dimensionality to produce a single output label. Here, the detection grid extends from $\hat{g}_{\text{min}}=0$ to $\hat{g}_{\text{max}}=1$, with a sampling resolution of $\delta\hat{g}=0.001$. The analysis employs a window size of $w=50$. The overall loss function landscape is constructed by integrating the losses from the trivial phase initialization for $g\leq 0.5$ with those from the topological phase initialization for $g>0.5$, for different values of $\delta$. Considering the notable volatility observed in the loss landscape, a Gaussian filter is applied to smooth the curve effectively. It is important to note that while the mathematical form of the order parameter learned here is not straightforwardly intuitive, it is encoded within the weights of the Transformer’s neural network layers. This order parameter can be readily calculated from measurement results using the recorded parameters of the Transformer, facilitating practical applications and analyses.
@@
The shaded areas in panel (d) of Fig. 7 reflect the phase diagram as reported in Ref. [22], which was calculated using the partial-reflection many-body topological invariant $\widetilde{\mathcal{Z}}_{\mathcal{R}}$ employing the infinite-size density matrix renormalization group technique [75]. We compute $\widetilde{\mathcal{Z}}_{\mathcal{R}}$ for both the optimized and true ground states of $\mathcal{H}_{\mathrm{eSSH}}$ at $n=16$. For $\delta<1.6$, the transition between the trivial and topological phases is marked where $\widetilde{\mathcal{Z}}_{\mathcal{R}}\approx 0$; for $\delta>1.6$, the transition between the trivial and symmetry-broken phases is identified where $\widetilde{\mathcal{Z}}_{\mathcal{R}}\approx 0.5$, and between the symmetry-broken and topological phases where $\widetilde{\mathcal{Z}}_{\mathcal{R}}\approx-0.5$. Green and blue hollow triangles represent the predicted phase boundaries from optimized and ground states, respectively, showing significant variance from those calculated using the partial-reflection invariant due to finite-size effects. Grey solid diamonds indicate the phase boundaries predicted by the Set Transformer for $\delta=0,0.5,1,1.5,\dots,4$, demonstrating a good agreement with the calculated ones. The small discrepancy is mainly due to residual finite-size effects.
@@
$\mathcal{H}_{\text{T-eSSH}}=\mathcal{H}_{\mathrm{eSSH}}(\delta,g)-h\sum_{j}^{n}\sigma^{x}_{j}.$ (35)
@@
Using exact diagonalization, we generate the ground states and corresponding feature vectors of $\mathcal{H}_{\text{T-eSSH}}$ for various $h$ and $g$ values, fixing $\delta=4$. We apply the recorded Transformer order parameters to learn the phase diagram, depicted in panel (e) of Fig. 7. Phase transitions are estimated at points where the Transformer model predicts a label of 0, indicative of shifts between phases. Notably, when $h<1.7$, the phase boundaries vary smoothly as a function of $h$; then, a sudden change occurs at $h=1.7$, and between $31.7\leq h<$ 3, the the boundaries are relatively stable as a function of $g$. Beyond $h\geq 3$, no stable phase transitions can be detected, indicating that the phase diagram is dominated by a single phase with respect to our order parameter in this regime.
IV Conclusions and Outlook
@@
In this work, we have presented a novel hybrid quantum optimization-machine learning algorithm designed to identify phase transitions in quantum systems. Through applications to the 1D and 2D transverse-field Ising models (TFIMs) and the extended Su-Schrieffer-Heeger (eSSH) model, we have demonstrated the algorithm’s robustness, versatility, and precision. Notably, our approach efficiently leverages locally optimized quantum states, obtained via low-depth variational quantum circuits, and classical machine learning techniques of LASSO and Transformer to predict critical points of phase transitions with remarkable accuracy.
Our findings illustrate that our algorithm is not only capable of detecting the critical points of phase transitions but also of unveiling novel order parameters. Notably, the LASSO algorithm identifies physically meaningful order parameters, thereby enhancing interpretability, whereas the Transformer uncovers complex, non-intuitive order parameters that, despite being less interpretable, indirectly provide crucial physical insights into the dynamics of phase transitions. This capability provides a novel tool to understand and analyse quantum phase transitions, extending beyond traditional methods that often rely on prior knowledge of suitable order parameters or are limited by computational scalability.
@@
Further enhancement of our algorithm could involve utilizing the differences in features rather than examining the features themselves. This strategy could enhance our understanding of transition dynamics, especially in systems where order parameters are not well-defined or universally applicable, such as in first-order phase transitions. Moreover, investigating our algorithm’s performance within the finite-temperature critical region surrounding quantum critical points represents a promising research direction. This exploration could provide deeper insights into how quantum fluctuations, thermal effects, and optimization deficits interact at criticality, thereby enriching our understanding of finite-temperature phase diagrams.

### v2

@@ II.3 Extended Su–Schrieffer–Heeger Model
In scenarios where the anisotropy parameter $\delta$ is small, increasing the coupling strength $g$ from 0 to 1 drives a phase transition from a trivial to a topological phase within the system [57]. Initially, the trivial phase displays a dimerized configuration, with spins forming singlet pairs predominantly with their nearest neighbours. As $g$ approaches 1, the system evolves into a topological phase, marked by the emergence of edge states that represent localized excitations at the system’s boundaries. Conversely, when $\delta$ exceeds a critical threshold (approximately $\delta^{\star}=1.6$), the system’s phase diagram becomes more complex, maintaining the trivial and topological phases while also exhibiting a symmetry-broken phase characterized by antiferromagnetic ordering. Notably, for values of $\delta\neq 0,1$, the model is non-integrable and lacks an analytical solution; hence, numerical and heuristic methods become essential for exploring its phase behavior.
@@
    $U^{\text{triv}}_{\text{eSSH}}(\boldsymbol{\beta},\boldsymbol{\gamma})=\prod_{j=1}^{p}$ $\exp\left(-i\frac{\beta_{j}}{2}\mathcal{H}_{xy}\right)\exp\left(-i\frac{\gamma_{j}}{2}\mathcal{H}_{zz}\right)$ (21)
@@
    $U^{\text{topo}}_{\text{eSSH}}(\boldsymbol{\beta},\boldsymbol{\gamma})=\prod_{j=1}^{p}$ $\exp\left(-i\frac{\beta^{\prime}_{j}}{2}\mathcal{H}^{\prime}_{xy}\right)\exp\left(-i\frac{\gamma^{\prime}_{j}}{2}\mathcal{H}^{\prime}_{zz}\right)$ (23)
@@
The optimization grid for the trivial phase initialization spans from $g_{\text{min}}=0$ to $g_{\text{max}}=0.6$, with a sampling resolution of $\delta g=0.05$. Conversely, the grid for the topological phase initialization ranges from $g_{\text{min}}=0.4$ to $g_{\text{max}}=1$, with the same sampling resolution of $\delta g=0.05$. Given the emphasis on symmetry-oriented phase transitions within this investigation, we first consider a scenario where $\mathcal{P}_{f}$ includes all non-identity, reflection-symmetric Pauli operators on subsystem $I$, detailed as follows:
@@
The simultaneous acquisition of their expectation values is facilitated through the execution of joint Bell measurements on pairs of qubits within $I$ that are symmetrically positioned, such as $(Q_{n/4},Q_{3n/4-1})$, $(Q_{n/4+1},Q_{3n/4-2})$, and so forth, up to $(Q_{n/2-1},Q_{n/2})$. Fig. 6(a) displays these joint Bell measurements performed on the corresponding qubit pairs in subsystems $I_{1}$ and $I_{2}$, illustrating the method for deriving the feature vector for a 16-qubit quantum chain. Notably, in the absence of noise, classically computing expectation values of such large-weight Pauli operators in $\mathcal{P}_{f}$ becomes significantly more resource-intensive, indicating a possible avenue towards practical quantum advantage.
@@
Figure 6: Overview of phase transition detection for the extended Su-Schrieffer-Heeger (eSSH) model. (a) Illustration of the eSSH model’s interaction scheme and the procedure for conducting joint Bell measurements on symmetrically positioned qubit pairs across subsystems $I_{1}$ and $I_{2}$ to derive the feature vector. (b) Fidelity comparison between the optimized quantum states and the true ground states, evaluated across various $g$ and $\delta$ values. (c) Plot of both the original and Gaussian-filter-smoothed loss landscapes for various $\delta$ values, underscoring the identification of critical points. Note that for $\delta=2,3,4$ two minima can be identified, signalling the presence of the two possible phase transitions between the trivial and topological phases, and between the trivial and symmetry-broken ones. (d) Comparison between the predicted phase boundaries, represented by green and blue hollow triangles (indicative of transitions computed from optimized and ground states using the partial-reflection many-body topological invariant $\widetilde{\mathcal{Z}}_{\mathcal{R}}$), and solid black diamonds representing the phase boundaries predicted by the Transformer model. This comparison is set against the phase diagram from Ref. [23], derived numerically using the infinite-size density matrix renormalization group technique. (e) Application of learned order parameters to the modified eSSH model with a transverse field ($\mathcal{H}_{\text{T-eSSH}}$), where $\delta=4$. This panel tracks the phase transitions as the field strength $h$ varies, with phase transitions identified where the Transformer model predicts shifts in phase labels, showcasing the adaptability of the learned parameters across different physical scenarios.
@@
To implement our algorithm, we employ a Set Transformer-based regressor, described in Section IV.2. This model is configured with four attention heads and operates at a learning rate of 0.001. Initially, the model applies a self-attention mechanism to the input features to discern dependencies and relationships. Following this, the architecture incorporates a predictive layer that performs a linear transformation, mapping the input features to an intermediate vector of dimension 128. This transformation is augmented by the ReLU activation function, which introduces non-linearity, and is followed by another linear transformation that reduces the feature dimensionality to produce a single output label. Here, the detection grid extends from $\hat{g}_{\text{min}}=0$ to $\hat{g}_{\text{max}}=1$, with a sampling resolution of $\delta\hat{g}=0.001$. The analysis employs a window size of $w=50$. The overall loss function landscape is constructed by integrating the losses from the trivial phase initialization for $g\leq 0.5$ with those from the topological phase initialization for $g>0.5$, for different values of $\delta$. Considering the notable volatility observed in the loss landscape, a Gaussian filter is applied to smooth the curve effectively. It is important to note that while the mathematical form of the order parameter learned here is not straightforwardly intuitive, it is encoded within the weights of the Transformer’s neural network layers. This order parameter can be readily calculated from measurement results using the recorded parameters of the Transformer, facilitating practical applications and analyses.
@@
The shaded areas in panel (d) of Fig. 6 reflect the phase diagram as reported in Ref. [23], which was calculated using the partial-reflection many-body topological invariant $\widetilde{\mathcal{Z}}_{\mathcal{R}}$ employing the infinite-size density matrix renormalization group technique [79]. We compute $\widetilde{\mathcal{Z}}_{\mathcal{R}}$ for both the optimized and true ground states of $\mathcal{H}_{\mathrm{eSSH}}$ at $n=16$. For $\delta<1.6$, the transition between the trivial and topological phases is marked where $\widetilde{\mathcal{Z}}_{\mathcal{R}}\approx 0$; for $\delta>1.6$, the transition between the trivial and symmetry-broken phases is identified where $\widetilde{\mathcal{Z}}_{\mathcal{R}}\approx 0.5$, and between the symmetry-broken and topological phases where $\widetilde{\mathcal{Z}}_{\mathcal{R}}\approx-0.5$. Green and blue hollow triangles represent the predicted phase boundaries from optimized and ground states, respectively, showing significant variance from those calculated using the partial-reflection invariant due to finite-size effects. Grey solid diamonds indicate the phase boundaries predicted by the Set Transformer for $\delta=0,0.5,1,1.5,\dots,4$, demonstrating a good agreement with the calculated ones. The small discrepancy is mainly due to residual finite-size effects. It is important to note that our method does not presuppose the number or location of phase transitions. Instead, the presence (or absence) of valleys in the loss landscape—observed during the variational optimization—serves as an indicator of phase transitions. Therefore, the tailored ansätze are employed solely to facilitate efficient convergence in the respective regimes, without relying on any prior knowledge of the eSSH model’s correct phase structure.
@@
$\mathcal{H}_{\text{T-eSSH}}=\mathcal{H}_{\mathrm{eSSH}}(\delta,g)-h\sum_{j}\sigma^{x}_{j}.$ (25)
@@
Using exact diagonalization, we generate the ground states and corresponding feature vectors of $\mathcal{H}_{\text{T-eSSH}}$ for various $h$ and $g$ values, fixing $\delta=4$. We apply the recorded Transformer order parameters to learn the phase diagram, depicted in panel (e) of Fig. 6. Phase transitions are estimated at points where the Transformer model predicts a label of 0, indicative of shifts between phases. Notably, when $h<1.7$, the phase boundaries vary smoothly as a function of $h$; then, a sudden change occurs at $h=1.7$, and between $31.7\leq h<$ 3, the boundaries are relatively stable with respect to $g$. For $h\geq 3$, no stable phase transitions can be detected, indicating that the phase diagram is dominated by a single phase with respect to our order parameter in this regime.
Next, we consider an alternative scenario in which no prior knowledge of the eSSH model is assumed. In this case, we construct $\mathcal{P}_{f}$ by including all local Pauli operators with weight two or less. Employing the same hyperparameters as described earlier, our methodology retains its ability to detect the critical phase boundaries and identify the emergence of the symmetry-broken phase. The Transformer loss landscapes are depicted in Fig. 7(a). They reveal two distinct minima for $\delta=2,3,4$, corresponding to the transitions between the trivial and symmetry-broken phases, and between the symmetry-broken and topological phases. The predictions remain largely consistent with the known phase boundaries, although finite-size effects become more pronounced in this setup, and the overall accuracy experiences a slight reduction, as shown in Fig. 7(b). These results suggest that the non-linear functions of 2-local Pauli terms learned by the Transformer model are sufficient to characterize the topological phase transitions in the eSSH model, even without explicit prior knowledge of the system’s properties.
Figure 7: (a) Plot of the original and Gaussian-filter-smoothed loss landscapes constructed by applying the Transformer model using expectation values of all local Pauli operators with weight below three for various $\delta$ values. For $\delta=2,3,4$, two minima are identified. (b) Comparison between the predicted phase boundaries, represented by solid black diamonds (Transformer model predictions from $2$-local Pauli terms), green hollow triangles (optimized states), and blue hollow triangles (ground states), against the phase diagram from Ref. [23], which was derived numerically using the infinite-size density matrix renormalization group technique.
III Discussion
@@
In this work, we have presented a hybrid quantum optimization-machine learning algorithm designed to identify phase transitions in quantum systems, leveraging locally trapped states produced by shallow variational circuits instead of exact ground-state data. This approach is particularly significant in the context of near-term quantum devices, where noise and circuit depth limitations preclude the accurate preparation of exact ground states. Through numerical simulations and real-hardware experiments of the 1D and 2D transverse-field Ising models (TFIMs) and the extended Su-Schrieffer-Heeger (eSSH) model, we have demonstrated the algorithm’s robustness and precision under the tested conditions.
Our real-hardware experiments validate the performance of our method under realistic noise conditions and represent, to our knowledge, one of the first experimental identifications of such order parameters directly from quantum data. Our findings illustrate that the algorithm not only detects the critical points of phase transitions but also unveils novel order parameters with faster convergence rates toward the thermodynamic limit and reduced sensitivity to finite-size effects in the systems investigated. Specifically, the LASSO algorithm identifies physically meaningful order parameters that enhance interpretability, while the Transformer model captures complex, non-intuitive many-body topological invariants that—although less directly interpretable—provide useful insights into phase transitions. Moreover, we have developed a new global optimization strategy that generates a continuous set of quantum states as a function of the Hamiltonian parameter $g$, thereby allowing precise tracking of phase evolution without requiring significantly deeper circuits.
We acknowledge that the system sizes investigated in this study are relatively modest, due to current hardware limitations and the need to measure numerous observables for training. While the results are encouraging, we are not claiming a formal quantum advantage here. Instead, the capacity of near-term devices to measure multiple observables in parallel suggests a potential practical speedup, especially if deeper or more adaptive measurements are employed. It is worth noting that while previous studies such as Herrmann et al. [80] and Cho et al. [81] have employed machine learning on experimental data, our approach addresses the more demanding task of identifying unknown quantum phase transitions using shallow-circuit variational optimization under realistic noise conditions.
An open question is the extent to which our approach—particularly in the shallow-circuit or local-observable regime—remains classically efficiently simulable [55, 82]. While shallow circuits with limited entanglement can, in some cases, be simulated using classical methods such as tensor networks [83, 84, 85], our algorithm is not inherently restricted to such settings. A central finding of our study is that precision improves exponentially with circuit depth, enabling accurate results with circuits of depth $O(\log n)$; notably, such circuits are not necessarily classically simulable. Finally, because our protocol does not require the ansatz to reach the exact ground state at every value of Hamiltonian parameter, we can use a curriculum approach: optimize the circuit in an easy, barren plateau‑free region [86], then transfer and fine‑tune the parameters as we move into classically intractable regimes.
Furthermore, being ‘classically simulable in principle’ rarely translates into practical feasibility. In many scenarios, the required algorithms suffer from large polynomial overheads or complex data structures that become prohibitive as $n$ grows [55]. Even if a shallow-circuit model is formally simulable, its resource demands may still outstrip available classical computing power for realistic system sizes. Mid-circuit measurements [87, 88] can further complicate classical strategies, and integrating our algorithm with established quantum-advantage protocols or fault-tolerant schemes [89, 90] may push shallow-depth training into regimes known to be classically intractable. While some configurations may appear classically tractable, the precise boundary between classical and quantum regimes remains unsettled, even in the shallow-depth, low-weight observable context explored here. Our results offer new insights into this ongoing debate.
@@
Future improvements to our algorithm could explicitly focus on analysing differences between feature sets, rather than individual features alone. This strategy could enhance our understanding of transition dynamics, especially in systems where order parameters are not well-defined or universally applicable, such as in first-order phase transitions. Moreover, investigating our algorithm’s performance within the finite-temperature critical region surrounding quantum critical points represents a promising research direction. This exploration could provide deeper insights into how quantum fluctuations, thermal effects, and optimization deficits interact at criticality, thereby enriching our understanding of finite-temperature phase diagrams.